# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ II Problem Formulation
Consider a population of $\mathsf{N}$ agents stratified into one anticoordinating and five coordinating subpopulations. The thresholds and population proportions of the subpopulations are respectively as follows: $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau^{\prime}_{5}},{\tau^{\prime}_{4}},{\tau^{\prime}_{3}},{\tau^{\prime}_{2}},{\tau^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.885},{0.89},{0.604},,{0.481},{0.444},{0.21})}$ $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho^{\prime}_{5}},{\rho^{\prime}_{4}},{\rho^{\prime}_{3}},{\rho^{\prime}_{2}},{\rho^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}12/30},{3/30},{3/30},{3/30},{1/30},{8/30}})$. For each value of $\mathsf{N}=30,60,120,\ldots,3840$, the population state was simulated 100 times with different random activation sequences and initial conditions. For each value of $\mathsf{N}$, we considered $30\mathsf{N}$ steps. For each run, the maximum and minimum of the population proportion of strategy $\mathtt{A}$ players were recorded. These values for each population size $\mathsf{N}$ are depicted in Figure 1. As population size grows, the maximum and minimum of the proportion of strategy-$\mathtt{A}$ players get closer and eventually merge.
Example 2 gives us an intuition that the amplitude of fluctuations in the population proportion of $\mathtt{A}$-players decreases as population size grows. Does this observation extend to every mixed population of coordinators and anticoordinators as the population size approaches infinity i.e., $\mathsf{N}\rightarrow\infty$? (Fig 2). The available results in stochastic approximation theory provide the basis for investigating this question.
Equilibrates [15]Equilibrates [16]May Fluctuate [19] Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both ? Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both $t\ \rightarrow\ \infty$$t\ \rightarrow\ \infty$ Population Population
Figure 2: How does the asymptotic behavior of the finite population of interacting agents (3) evolve as the population size approaches infinity? An exclusive finite population of coordinators and an exclusive finite population of anticoordinators each equilibrates in the long-run, [15, 16]. In the latter case, for some specific configurations, the population may admit a two-state cycle. The proportion of $\mathtt{A}$-players in a mixed finite population may undergo perpetual fluctuations [19].
@@ III Background
where $\bm{\mathcal{V}}\subset\mathbb{R}^{n}$ is a set-valued map specifying a set of evolution at each point $\bm{x}\in\mathbb{R}^{n}$.
Definition 3 ([38]).
@@
Each $\bm{y}$ in Definition 3 is a solution of the differential inclusion, and the term $\bigcap_{t\geq 0}\text{cl}(\bm{y}[t,\infty])$ is the limit set of $\bm{y}$. In words, $\bm{x}_{0}$ is a recurrent point of $\bm{\Phi}$ if it is included in the limit set of some solution that starts from $\bm{x}_{0}$. Namely, a recurrent point generalizes the notion of an equilibrium point as, for example, the points on a limit cycle are recurrent points.
@@
[38] Consider the good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ over the convex and compact state space $\bm{\mathcal{X}}_{0}$. For a sequence of positive values of scalar $\epsilon$ approaching 00, let $\textbf{\em U}^{\epsilon}=\langle\textbf{\em U}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ be a sequence of $\mathbb{R}^{n}$-valued random variables and $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ be a family of set-valued maps on $\mathbb{R}^{n}$. We say that $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\epsilon>0}$ is a family of generalized stochastic approximation processes (or a GSAP) for the differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ if the following conditions are satisfied:
@@
    $\mathbf{X}^{\epsilon}_{k+1}-\mathbf{X}^{\epsilon}_{k}-\epsilon\textbf{\em U}^{\epsilon}_{k+1}\in\epsilon\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$,
@@
    $\lim_{\epsilon\rightarrow 0}\mathbb{P}\left[\max_{k\leq\frac{T}{\epsilon}}\left|\textstyle\sum_{i=1}^{k}\epsilon\textbf{\em U}^{\epsilon}_{i}\right|>\alpha\mid\mathbf{X}^{\epsilon}_{0}=\bm{x}\right]=0$
@@
The GSAPs can be regarded as perturbed solutions of the differential inclusions. The first condition ensures that the perturbed solutions are within the same set as the unperturbed solutions are. The second condition defines $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ as a sequence of elements which are recursively updated. At each index $k$, the value of $\mathbf{X}^{\epsilon}_{k+1}$ is characterized based on those of $\mathbf{X}^{\epsilon}_{k}$, $\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$, and the random variable $\textbf{\em U}^{\epsilon}_{k+1}$. The third condition ensures that the collection $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ approaches to the vicinity of $\bm{\mathcal{V}}$ as $\epsilon$ approaches zero. More specifically, the third condition is indeed a limit condition, which requires that for an arbitrarily small $\delta>0$ there exists $\epsilon_{0}>0$ such that for each $\epsilon<\epsilon_{0}$ in the sequence $\langle\epsilon\rangle$, the values belonging to the set valued map $\bm{\mathcal{V}}^{\epsilon}(\bm{x})$ are within a $\delta$-neighborhood of the image of the $\delta$-neighborhood of $\bm{x}$. Well-behaved random variables, such as sub-Gaussian random variables and some classes of random variables with finite higher moments satisfy the last condition.
In [38], Roth and Sandholm studied the limiting stationary distributions of these GSAPs and showed that if for each $\epsilon>0$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle$ is a Markov chain, their stationary measures as $\epsilon$ approaches zero will be concentrated on $\mathrm{BC}(\bm{\Phi})$. The following theorem is a result of A1 and was implicitly used in the proof of [38, Theorem 4.11].
@@ IV The link to the continuous-time dynamics
In view of (3), the next state of the population dynamics only depends on the current population state and the adopted strategy of and the subpopulation associated with the current active agent. As a result, the sequence of $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{{}_{k=0}^{\infty}}}$ is a realization of a Markov chain.
@@
$\begin{cases}({\rho}_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))&\text{if }\bm{y}^{\mathsf{N}}=\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=-\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ {1-}\displaystyle\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\big{(}x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=\bm{x}^{\mathsf{N}},\\ \hskip 27.0pt+(\rho_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))\big{)}\\ 0&\text{otherwise},\end{cases}$ (6)
@@
For the sake of readability, the proofs of the results are provided in the appendix.
@@
We increase the population size $\mathsf{N}$ in a way that the population structure, i.e., the population proportions of the subpopulations $\bm{\rho}$ remains unchanged. Hence, the sequence according to which the population size approaches infinity should satisfy $\mathsf{N}\bm{\rho}\in\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$. We denote the set of such valid population sizes by $\mathcal{N}$.
Now, to apply Theorem 1, we need to show that the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for a good upper semicontinuous differential inclusion and hence the support of the limit point of its invariant probability measures is determined by the steady-state behavior of that differential inclusion. We claim that $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the following differential inclusion:
@@
${\mathcal{V}}_{p}(\bm{x})=\begin{cases}\{{\rho}_{p}-x_{p}\}&\text{if }x<\tau_{p},\\ [-x_{p},{\rho}_{p}-x_{p}]&\text{if }x=\tau_{p},\\ \{-x_{p}\}&\text{if }x>\tau_{p},\end{cases}$ (7a)
@@
The above dynamics are indeed the mean dynamics of the discrete population dynamics (3) [20]. The mean dynamic considers the decision-making process as a continuous process with an infinitely large population, and $\dot{x}_{p}$ describes the rate of change in the population proportion of strategy-$\mathtt{A}$ players in subpopulation $p$. This equals the expected inflow,$F_{p,\mathtt{A}}$, that is, the rate of change to strategy $\mathtt{A}$ minus the expected outflow, $F_{p,\mathtt{B}}$, that is, the rate of change to strategy $\mathtt{B}$:
$\dot{x}_{p}\in{\{}F_{p,\mathtt{A}}(\bm{x})-F_{p,\mathtt{B}}(\bm{x}){\}},$ (8)
where $F_{p,\mathtt{A}},F_{p,\mathtt{B}}:\bm{\mathcal{X}}_{ss}\rightarrow{[0,\rho_{p}]}$. If the population proportion of strategy-$\mathtt{A}$ players is greater (resp. less) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the inflow will be zero, and the outflow equals the current proportion of strategy-$\mathtt{A}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=-x_{p}$. On the contrary, if the population proportion of strategy-$\mathtt{A}$ players is less (resp. greater) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the outflow will be zero and the inflow equals the current proportion of strategy $\mathtt{B}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=\rho_{p}-x_{p}$. When the population proportion of strategy-$\mathtt{A}$ players equals the threshold of (anticoordinating or coordinating) subpopulation $p$, both strategies $\mathtt{A}$ and $\mathtt{B}$ are a legitimate choice should there be no tie-breaker, and some may choose strategy $\mathtt{A}$ while others may choose strategy $\mathtt{B}$. Therefore, the rate will be bounded by the above two cases, i.e., $-x_{p}$ and $\rho_{p}-x_{p}$, resulting in the interval $[-x_{p},\rho_{p}-x_{p}]$. If there is a tie-breaker, such as the set-up in this paper, then one of the strategies is preferred, resulting in one of the above two cases, which results in a selection from the defined differential inclusion. A selection from $\bm{\mathcal{V}}(\bm{x})$ is a singleton map $\bm{\nu}(\bm{x})\in\mathbb{R}^{n}$ which satisfies $\bm{\nu}(\bm{x})\in\bm{\mathcal{V}}(\bm{x})$ for all $\bm{x}$ in $\mathbb{R}^{n}$ [42]. That is why the mean dynamics of the discrete population dynamics (3) is sometimes considered as a selection from (7) [38]. It can be easily checked that the differential inclusion (7) satisfies the conditions of good upper semicontinuity.
@@
The collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for (7).
Based on Lemma 1, the population dynamics (3) define GSAPs for the differential inclusion (7). Hence, in view of Theorem 1 the support of the limit point of the stationary distributions of (6), as $\mathsf{N}\rightarrow\infty$, is contained in the Birkhoff center of $\bm{\Phi}$ induced by (7). In the next section, we determine the Birkhoff center of $\bm{\Phi}$.
@@
The set-up in (1), (2), and (3) is based on a tie-breaking rule [43] where the agents prefer strategy $\mathtt{A}$ when the population proportion of $\mathtt{A}$-players, excluding themselves, equals the agent’s threshold, i.e., $x^{\mathsf{N}}_{-i}=\tau(i)$ for every agent $i$. However, for other tie-breaking rules, such as when strategy $\mathtt{B}$ is preferred, or the agents uniformly randomly choose either of the strategies, Lemma 1 remains valid–A1. The reason is that a different tie-breaker will only change $s^{*}(\cdot,\cdot,\cdot)$ in Definition 1 and consequently the selection from (7) when $x=\tau_{p}$ where $p$ is the type of agent $i$. Accordingly, the dynamical behavior of the resulting selection is already captured by (7). Therefore, investigating the asymptotic behavior of (7) does reveal that of the discrete population dynamics (3) with such tie-breaking rules as population size approaches infinity.

### v4

@@ II Problem Formulation
Consider a population of $\mathsf{N}$ agents stratified into one anticoordinating and five coordinating subpopulations. The thresholds and population proportions of the subpopulations are, respectively, as follows: $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau^{\prime}_{5}},{\tau^{\prime}_{4}},{\tau^{\prime}_{3}},{\tau^{\prime}_{2}},{\tau^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.885},{0.89},{0.604},,{0.481},{0.444},{0.21})}$ $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho^{\prime}_{5}},{\rho^{\prime}_{4}},{\rho^{\prime}_{3}},{\rho^{\prime}_{2}},{\rho^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}12/30},{3/30},{3/30},{3/30},{1/30},{8/30}})$. For each value of $\mathsf{N}=30,60,120,\ldots,3840$, the population state was simulated 100 times with different random activation sequences and initial conditions. For each value of $\mathsf{N}$, we considered $30\times\mathsf{N}$ steps. For each run, the maximum and minimum of the population proportion of strategy $\mathtt{A}$ players over the last $2\mathsf{N}$ steps were recorded. These values for each population size $\mathsf{N}$ are depicted in Figure 1. As the population size grows, the maximum and minimum of the proportion of $\mathtt{A}$-players get closer.
Example 2 gives us an intuition that the amplitude of fluctuations in the population proportion of $\mathtt{A}$-players decreases as the population size grows. Does this observation extend to every mixed population of coordinators and anticoordinators as the population size approaches infinity i.e., $\mathsf{N}\rightarrow\infty$? (Fig 2). The available results in stochastic approximation theory provide the basis for investigating this question.
Equilibrates [ramazi2020convergence]Equilibrates [ramazi2017asynchronous]May Fluctuate [roohi] Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both ? Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both tt\ $\rightarrow\ \inftytt\ \rightarrow\ \infty$ Population Population
Figure 2: How does the asymptotic behavior of the finite population of interacting agents (3) evolve as the population size approaches infinity? A finite population of all coordinators and a finite population of all anticoordinators each equilibrates in the long run, [ramazi2020convergence, ramazi2017asynchronous]. In the latter case, for some specific configurations, the population may admit a two-state cycle. The proportion of $\mathtt{A}$-players in a mixed finite population may undergo perpetual fluctuations [roohi].
@@ III Background
where $\bm{\mathcal{V}}$ is a set-valued map specifying a set of evolutions in $\mathbb{R}^{n}$ at each point $\bm{x}\in\mathbb{R}^{n}$. A Caratheodory solution of (4) on $0,t_{1}]\subset\mathbb{R}_{\geq 0}$ is an absolutely continuous map $\bm{x}:[0,t_{1}]\rightarrow\mathbb{R}^{n}$ such that for almost all $t\in[0,t_{1}]$, $\dot{\bm{x}}(t)\in\bm{\mathcal{V}}(\bm{x}(t))$ [cortes2008discontinuous]. Robotic manipulators interacting with objects, high-speed supercavitation vehicles, and power systems can be well modeled and analyzed within the framework of differential inclusions [cortes2008discontinuous, kani, powerSystem]. A point $\bm{x}^{*}\in\mathbb{R}^{n}$ that satisfies ${\mathbf{0}}\in\bm{\mathcal{V}}(\bm{x}^{*})$ is an equilibrium of (4) [cortes2008discontinuous]. The set-valued dynamical system induced by the differential inclusion (4) is denoted by $\bm{\Phi}(\bm{x}_{0})$ and defined as the set of all solutions of (4) with initial condition $\bm{x}_{0}$ [roth2013stochastic]. A set-valued map $\bm{\mathcal{V}}$ is called upper semicontinuous if its graph, $\{(\bm{x},\bm{y})|\bm{y}\in\bm{\mathcal{V}}(\bm{x})\}$, is closed. When an upper semicontinuous map is equipped with three additional properties of nonemptiness, convexity, and boundedness, the map is good upper semicontinuous.
Definition 3 ([roth2013stochastic]).
@@
Each $\bm{y}$ in Definition 3 is a solution of the differential inclusion, and the term $\bigcap_{t\geq 0}\text{cl}(\bm{y}[t,\infty))$ is the limit set of $\bm{y}$. In words, $\bm{x}_{0}$ is a recurrent point of $\bm{\Phi}$ if it is included in the limit set of some solution that starts from $\bm{x}_{0}$. Namely, a recurrent point generalizes the notion of an equilibrium point, as, for example, the points on a limit cycle are recurrent.
@@
[roth2013stochastic] Consider the good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ over the convex and compact state space $\bm{\mathcal{X}}_{0}$. For a sequence of positive values of scalar $\epsilon$ approaching 0, let $\textbf{U}^{\epsilon}=\langle\textbf{U}^{\epsilon}_{k}\rangle{{}_{k=0}^{\infty}}$ be a sequence of $\mathbb{R}^{n}$-valued random variables and $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ be a family of set-valued maps on $\mathbb{R}^{n}$. We say that $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{}_{k=0}^{\infty}}\rangle_{\epsilon>0}$ is a family of generalized stochastic approximation processes (or a GSAP) for the differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ if the following conditions are satisfied:
@@
    $\mathbf{X}^{\epsilon}_{k+1}-\mathbf{X}^{\epsilon}_{k}-\epsilon\textbf{U}^{\epsilon}_{k+1}\in\epsilon\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$,
@@
    $\lim_{\epsilon\rightarrow 0}\mathbb{P}\left[\max_{k\leq\frac{T}{\epsilon}}\left|\textstyle\sum_{i=1}^{k}\epsilon\textbf{U}^{\epsilon}_{i}\right|>\alpha\mid\mathbf{X}^{\epsilon}_{0}=\bm{x}\right]=0$
@@
The GSAPs can be regarded as perturbed solutions of the differential inclusions. The first condition ensures that the perturbed solutions remain in the same set as the unperturbed ones. The second condition defines $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{}_{k=0}^{\infty}}$ as a sequence of elements which are recursively updated. At each index $k$, the value of $\mathbf{X}^{\epsilon}_{k+1}$ is characterized based on those of $\mathbf{X}^{\epsilon}_{k}$, $\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$, and the random variable $\textbf{U}^{\epsilon}_{k+1}$. The third condition ensures that the collection $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ approaches to the vicinity of $\bm{\mathcal{V}}$ as $\epsilon$ approaches zero. More specifically, the third condition is indeed a limit condition, which requires that for an arbitrarily small $\delta>0$ there exists $\epsilon_{0}>0$ such that for each $\epsilon<\epsilon_{0}$ in the sequence $\langle\epsilon\rangle$, the values belonging to the set valued map $\bm{\mathcal{V}}^{\epsilon}(\bm{x})$ are within a $\delta$-neighborhood of the image of the $\delta$-neighborhood of $\bm{x}$. Well-behaved random variables, such as sub-Gaussian random variables and some classes of random variables with finite higher moments, satisfy the last condition.
In [roth2013stochastic], Roth and Sandholm studied the limiting stationary distributions of these GSAPs and showed that if for each $\epsilon>0$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle$ is a Markov chain, their stationary measures as $\epsilon$ approaches zero will be concentrated on $\mathrm{BC}(\bm{\Phi})$. The following theorem is a result of A1 and was implicitly used in the proof of [roth2013stochastic, Theorem 4.11].
@@ IV The link to the continuous-time dynamics
In view of (3), the next state of the population dynamics only depends on the current population state, the strategy of and the subpopulation associated with the current active agent. As a result, the sequence of $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{}_{k=0}^{\infty}}$ is a realization of a Markov chain.
@@
$\begin{cases}({\rho}_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},2))&\hskip-35.0pt\text{if }\bm{y}^{\mathsf{N}}=\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},1)-1)&\hskip-35.0pt\text{if }\bm{y}^{\mathsf{N}}=-\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ 1-\displaystyle\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\big(x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},1)-1)\hskip-2210.0pt&\hskip-35.0pt\text{if }\bm{y}^{\mathsf{N}}=\bm{x}^{\mathsf{N}},\\ \hskip 27.0pt+(\rho_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},2))\big)\\ 0&\hskip-35.0pt\text{otherwise},\end{cases}$ (6)
@@
See Appendix for a proof.
@@
We increase the population size $\mathsf{N}$ in a way that the population structure, i.e., the population proportions of the subpopulations (vector $\bm{\rho}$) remains unchanged. Hence, the sequence according to which the population size approaches infinity should satisfy $\mathsf{N}\bm{\rho}\in\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$. We denote the set of such valid population sizes by $\mathcal{N}$.
Now, to apply Theorem 1, we need to show that the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{}_{k=0}^{\infty}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for a good upper semicontinuous differential inclusion, and hence the support of the limit point of its invariant probability measures is determined by the steady-state behavior of that differential inclusion. We claim that $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{}_{k=0}^{\infty}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the following differential inclusion:
@@
$\hskip-55.0pt{\mathcal{V}}_{p}(\bm{x})=\begin{cases}\{{\rho}_{p}-x_{p}\}&\text{if }x<\tau_{p},\\ [-x_{p},{\rho}_{p}-x_{p}]&\text{if }x=\tau_{p},\\ \{-x_{p}\}&\text{if }x>\tau_{p},\end{cases}$ (7a)
@@
The above dynamics are indeed the mean dynamics of the discrete population dynamics (3) [sandholm2010population]. The mean dynamic considers the decision-making process as a continuous process with an infinitely large population, and $\dot{x}_{p}$ describes the rate of change in the population proportion of strategy-$\mathtt{A}$ players in subpopulation $p$. This equals the expected inflow, $F_{p,\mathtt{A}}$, that is, the rate of change to strategy $\mathtt{A}$ minus the expected outflow, $F_{p,\mathtt{B}}$, that is, the rate of change to strategy $\mathtt{B}$:
$\dot{x}_{p}\in\{a-b|a\in F_{p,\mathtt{A}}(\bm{x}),b\in F_{p,\mathtt{B}}(\bm{x})\},$ (8)
where $F_{p,\mathtt{A}},F_{p,\mathtt{B}}:\bm{\mathcal{X}}_{ss}\rightrightarrows{[0,\rho_{p}]}$. If the population proportion of strategy-$\mathtt{A}$ players is greater (resp. less) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the inflow will be zero, and the outflow equals the current proportion of strategy-$\mathtt{A}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=-x_{p}$. On the contrary, if the population proportion of strategy-$\mathtt{A}$ players is less (resp. greater) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the outflow will be zero and the inflow equals the current proportion of strategy $\mathtt{B}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=\rho_{p}-x_{p}$. When the population proportion of strategy-$\mathtt{A}$ players equals the threshold of (anticoordinating or coordinating) subpopulation $p$, both strategies $\mathtt{A}$ and $\mathtt{B}$ are a legitimate choice should there be no tie-breaker, and some may choose strategy $\mathtt{A}$ while others may choose strategy $\mathtt{B}$. Therefore, the rate will be bounded by the above two cases, i.e., $x_{p}$ and $\rho_{p}-x_{p}$, resulting in the interval $x_{p},\rho_{p}-x_{p}]$. If there is a tie-breaker, such as the set-up in this paper, then one of the strategies is preferred, resulting in one of the above two cases, which results in a selection from the defined differential inclusion. A selection from $\bm{\mathcal{V}}(\bm{x})$ is a singleton map $\bm{\nu}(\bm{x})\in\mathbb{R}^{n}$ which satisfies $\bm{\nu}(\bm{x})\in\bm{\mathcal{V}}(\bm{x})$ for all $\bm{x}$ in $\mathbb{R}^{n}$ [smirnov2002introduction]. That is why the mean dynamics of the discrete population dynamics (3) is sometimes considered as a selection from (7) [roth2013stochastic]. By extending the domain of the continuous-time population dynamics to $\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$ using projection and in view of [sandholm2010population, Theorem 6.A.2], the good upper semicontinuity of the differential inclusion (7) can be verified.
@@
Assume that the thresholds of the subpopulations are distinct. Then the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{}_{k=0}^{\infty}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a family of GSAPs for (7).
For the sake of readability, the proofs of the results are provided in the appendix. Based on Lemma 1, the population dynamics (3) define GSAPs for the differential inclusion (7). Hence, in view of Theorem 1, the support of the limit point of the stationary distributions of (6), as $\mathsf{N}\rightarrow\infty$, is contained in the Birkhoff center of $\bm{\Phi}$ induced by (7). In the next section, we determine the Birkhoff center of $\bm{\Phi}$.
@@
The set-up in (1), (2), and (3) is based on a tie-breaking rule [sakhaei2022equilibration] where the agents prefer strategy $\mathtt{A}$ when the population proportion of $\mathtt{A}$-players, excluding themselves, equals the agent’s threshold, i.e., $x^{\mathsf{N}}_{-i}=\tau(i)$ for every agent $i$. However, for other tie-breaking rules, such as when strategy $\mathtt{B}$ is preferred, or the agents uniformly randomly choose either of the strategies, Lemma 1 remains valid (see Corollary A1). The reason is that a different tie-breaker will only change $s^{*}(\cdot,\cdot,\cdot)$ in Definition 1 and consequently the selection from (7) when $x=\tau_{p}$ where $p$ is the type of agent $i$. Accordingly, the dynamical behavior of the resulting selection is already captured by (7). Therefore, investigating the asymptotic behavior of (7) does reveal that of the discrete population dynamics (3) with such tie-breaking rules as population size approaches infinity.